# Redshift Assessment Infrastructure Layers (RAIL): Rubin-era photometric redshift stress-testing and at-scale production

### v2

@@ 1 Introduction
The Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) will obtain deep $ugrizy$ imaging of $\sim 18,000$ square degrees optimized for extragalactic science during its ten-year survey. It will produce a catalog of 20 billion galaxies down to $i=26.4$, enabling advances in our understanding of numerous extragalactic phenomena and cosmology (Ivezić et al., 2019). Most, if not all, such studies require some notion of the distance or redshift of each galaxy, and in some cases the distribution of redshifts for ensembles of galaxies (see Table 1 of Breivik et al. 2022). Redshift may be measured directly from the absorption and emission lines in a spectrum. However, LSST’s galaxy sample poses a twofold challenge to redshifts measured directly from the absorption and emission lines in a spectrum: the large number of galaxies greatly exceeds the available time on spectroscopic follow-up observing facilities, and the faint sources that comprise the bulk of the sample will be inaccessible to spectroscopic instruments regardless.
In lieu of spectroscopically confirmed redshifts, LSST will yield photometric redshifts (photo-$z$s) derived from broadband photometry, a standard data product of photometric galaxy surveys (e.g., Tanaka et al., 2018; Bilicki et al., 2018; Buchs et al., 2019). Photo-$z$s are subject to a variety of sources of inaccuracy and imprecision that are anticipated to be more severe for LSST’s faint sample, making point estimates and Gaussian uncertainties inappropriate approximations for nearly all extragalactic applications of LSST data. Photo-$z$ probability density functions (PDFs) comprehensively characterize the redshift uncertainty (Tanaka et al., 2018; Bilicki et al., 2018; Buchs et al., 2019) and are thus favored for LSST data. A thorough review of photo-$z$ uncertainty characterization can be found in Newman & Gruen (2022).
@@ 2.3 Data Handle
The tabular data handles (TableHandle) wrap catalog-like data, which can exist as a Numpy dictionary, pandas dataframe, pyarrow table, or Astropy table. The TableHandle interacts with tables_io to handle different types of file formats. More information can be found about tables_io in Appendix B.1. The qp handle (QPHandle) wraps qp ensembles (Malz & Marshall, 2017), which are iterable data structures of 1D PDFs. Additional information about qp can be found in Appendix B.2. The model handle (ModelHandle) wraps the model generated by a particular algorithm in RAIL, informed by some training data. Regardless of their types, the models are stored in pickle files by RAIL.
@@ 2.4 Definitions
Photo-$z$ probability density distribution (PDF) $p(z)$ A photo-$z$ PDF in this paper is referred to the posterior distribution of the redshift of an individual galaxy. The conditions (i.e., given the galaxy’s photometry) is typically omitted, but should not be forgotten.
@@
Photometric data $\mathbf{p}$ Photometric data in this paper is referred to the galaxy catalog information containing photometry (i.e. fluxes or magnitudes in different filters). This can refer to both simulated and real data.
@@
Color $c$ A color is defined as $c=m_{I}-m_{I^{\prime}}$, where $I$ and $I^{\prime}$ are adjacent filters. It is also often referred to as $I-I^{\prime}$. For example, $u-g$ or $g-i$ color.
@@
True redshift PDF $p(z_{t}|\mathbf{p}_{t})$ The conditional PDF of the true redshift, given the true photometry of a galaxy. In RAIL this is provided for galaxies simulated using a probabilistic model for which the joint distribution of redshift and magnitudes from which the galaxies are sampled is known exactly. For example, this is true of galaxies simulated using PZFlow. In Section 5, we also refer to this as $p_{\rm true}(z)$.
@@
Figure 2: The workflow of the RAIL.creation forward modeling subpackage. Input and output data are represented by rectangles, and RAIL stages are represented by ovals. A typical creation pipeline starts with training a creation (‘generative’) model from either a reference catalog (often simulations), such as in the case of PZFlow, or from the template SEDs drawn from a prior, such as in the case of DSPS/FSPS. The empirical or explicit joint distribution of true redshifts with other catalog properties (‘data’) is then computed. A mock truth catalog (‘underlying galaxy catalog’) is then drawn from the creation model. This catalog is then degraded by the degradation stages, such that it mimics a noisy observed catalog (‘biased galaxy sample’).
@@ 3 Creation and Degradation
Mock DESC data are important for systematically testing the performance of various photo-$z$ algorithms. One of the lessons learned from DC1 is that it is desirable for the mock data to include not only true redshifts and LSST photometry (i.e., fluxes in the six LSST bands) but also true posterior PDFs, $p(z_{t}|\mathbf{p}_{t})$, which are unavailable for spectroscopically confirmed data sets as well as traditional simulations. Furthermore, the mock photometry data should contain realistic noise, selection effects, and biases. This is critical for the training and validation of photo-$z$ algorithms.
@@ 3.1 Engines
An engine is defined by a pair of stages that are subclasses of each of the following superclasses: RAIL.creation.Modeler makes a model of the $p(z,\mathrm{photometry})$ joint probability space based on input parameters or data, and RAIL.creation.Creator samples $(z,\mathrm{photometry})$ from the forward model. Available engines are listed in Table 2.
@@
FSPS Physical rail-fsps Conroy et al. (2009); Conroy & Gunn (2010)
@@ 3.1.1 FSPS (Flexible Stellar Population Synthesis)
FSPS is a RAIL module that creates an interface to the Python bindings of the popular stellar population synthesis (SPS) code FSPS (Flexible Stellar Population Synthesis, Conroy et al. 2009; Conroy & Gunn 2010). FSPS aims at generating realistic galaxy spectral energy distributions (SEDs) by modelling all the components that contribute to the light from a galaxy: stars, gas, dust and AGN. FSPS is widely used both for stellar population inference (Johnson et al., 2021) and for forward modelling of galaxy SEDs (e.g., Alsing et al. 2023; Tortorelli et al. 2024).
@@ 3.2.4 QuantityCut
This degrader provides a trimmed version of the input catalog based on selection cuts applied to the catalog quantities. The user provides the parameter cuts, which is a dictionary with keys being the columns to which the selection is to be applied (e.g., the $i$-band magnitude), and the values being the specific cuts. Two types of values can be provided: a single float number (e.g., 25.3), which is interpreted as a maximum value (i.e., the cut will remove samples with $i>25.3$), and a tuple (e.g., $(17,25.3)$), which is interpreted as a range within which the sample is selected (i.e., the selected sample has $25.317<i<$ 25.3). When multiple cuts are applied at the same time, only the intersection of selected samples of each cut will be kept in the output.
@@ 3.2.7 Blending Degrader
[…]
@@ 4 Photo-$z$ Estimation
Delight CatInformer, CatEstimator rail-delight Leistedt & Hogg (2017)
@@
FlexZBoost CatInformer, CatEstimator rail-flexzboost Izbicki & Lee (2017b)
@@
TPZ CatInformer, CatEstimator rail-tpz Carrasco Kind & Brunner (2013)
@@
Naive stacking PzInformer, PZSummarizer rail-base Malz & Hogg (2020)
@@ 4.1.3 FlexZBoost
FlexZBoost (Izbicki & Lee, 2017b; Dalmasso et al., 2020) is an algorithm based on conditional density estimate that uses the FlexCode 777available at https://github.com/lee-group-cmu/FlexCode package. The package parameterises the PDF as a linear combination of orthonormal basis functions (a set of unit vectors in the color space that are orthogonal to each other), where the basis function coefficients can be determined by regression. The RAIL implementation uses xgboost (Chen & Guestrin, 2016) to perform the regression. The basis function representation of the photo-z PDF of a galaxy can lead to small-scale residual ‘bumps’. In the course of training the density estimate, an optimal threshold (configuration parameter bump_thresh) below which small-scale features are removed is determined by setting aside a fraction of the training data and minimizing the CDE loss at different threshold values. Additionally, the width of the final PDF is similarly optimized by the inclusion of a ‘sharpening’ parameter that scales the PDF by a power law value $\alpha$. Again a fraction of the training data is set aside and the CDE loss is minimized over a set of $\alpha$ values. The resultant photo-$z$ PDF distributions can be stored as qp.Ensembles either in their native basis function representation, or as a linearly interpolated grid.
@@ 4.1.4 GPz
GPz is an algorithm based on sparse Gaussian Process, which was introduced by Almosallam et al. (2016). The current RAIL implementation of GPz is the preliminary version; that is, it predicts a single Gaussian PDF rather than the more sophisticated multimodal PDFs that are implemented in newer versions of GPz Stylianou et al. (2022). GPz models both the mean and standard deviation of the Gaussian PDF as a linear combination of basis functions, and learns the parameters for the basis functions via a Gaussian process. The method can make one of several assumptions about the covariance between these basis functions, which are controlled via the configuration parameter gpz_method as outlined in the RAIL documentation.
@@ 4.2.1 BPZ (Bayesian Photometric Redshifts)
BPZ is a template-based estimator developed by Benítez (2000). Like many template-based codes, it operates by computing synthetic fluxes for an input set of SEDs by integrating the products of the SEDs and the filter bandpass curves for a particular survey. For each SED the algorithm computes a $\chi^{2}$ value using the (scaled) fluxes and the observed data via:
$\chi^{2}=\sum_{\alpha}\frac{(f_{\alpha}-a_{0}f_{T\alpha})^{2}}{2\sigma_{f\alpha}^{2}}$ (4)
where the summation $\sum_{\alpha}$ is over the number of available filters, $f_{\alpha}$ is the observed flux in band $\alpha$, $a_{0}$ is an arbitrary scaling factor (the use of this scaling factor means that the template SEDs have constant shape and thus do not evolve as a function of luminosity), $f_{T\alpha}$ is the model flux of template $T$ in band $\alpha$, and $\sigma_{f\alpha}$ is the observed flux uncertainty in band $\alpha$. These $\chi^{2}$ values are converted into likelihoods, and an SED-type-dependent apparent magnitude prior is applied to the likelihoods before they are marginalized over type to produce a final posterior probability distribution.
The BPZliteEstimator stage takes a TableHandle catalog of magnitudes and magnitude errors as input, and returns an interpolated grid qp.Ensemble of posterior PDFs. As the likelihood values are computed on a grid, the mode values for each galaxy as measured on the grid are also returned by default. Also included in the ancillary data are values tb corresponding to the ‘best-fit SED type’ (evaluated at the mode redshift), and todds, a parameter that gives the fraction of the probability that comes from SED type tb at the mode redshift. Low values of todds mean that multiple SEDs are contributing to the probability total at the mode redshift, and thus a ‘best fit type’ is ill-defined, while values close to unity mean that most or all of the probability is from a single SED type, and thus the use of a ‘best fit type’ may be appropriate for the individual galaxy.
BPZ adopts a prior for redshifts and galaxy types: $\pi(z,T|m_{0})=\pi(z|T,m_{0})\pi(T|m_{0})$, where $\pi(T|m_{0})$ is the spectral type prior (i.e., the galaxy type fraction as a function of magnitude), and $\pi(z|T,m_{0})$ is the redshift prior. The function forms of these priors are given by Eq. 28 - 30 in Benítez (2000), with a total of 11 free parameters. A novel feature introduced in rail_bpz and not available in the original code is the ability for the user to train a custom apparent magnitude prior, which is accessed via the BPZliteInformer stage. However, the user must first classify the training data into SED ‘types’ before this feature is run. This step is not necessary, and rail_bpz can be run without training using the ‘default’ prior that is used in the original code. There is no guarantee that the returned prior will be robust, given the complex nature of potential input data configurations. Therefore, this feature should be used with caution, and users should check that the resulting priors are sensible before incorporating them in any analysis.
@@ 4.3.1 Delight
Leistedt & Hogg (2017) introduced a novel approach to inferring photometric redshifts which combines some of the strengths of machine learning and template-fitting methods by implicitly constructing flexible template SEDs directly from the spectroscopic training data, called Delight. It is a method for calculating the posterior probability of redshift given a catalog of deep observations acting as a data-driven prior. The catalog can have observations in arbitrary bands and with arbitrary noise; Gaussian processes are used as a principled method to implicitly construct SEDs (capturing the effects of redshifts, bandpasses and noise). The hyperparameters of the Gaussian process can be optimized as a calibration step.
@@ 4.5.3 Naive methods
The NaiveStackSummarizer takes the photo-$z$ PDF from each catalog object and simply average them over the ensemble to produce the ensemble redshift distribution $n(z)$, i.e., $n(z)=\sum_{i}^{N}p_{i}(z)/N$. A set of bootstrap realizations is also produced to estimate the uncertainties on the distributions. The NaiveStackMaskedSummarizer can deal with tomographic bins. Given the tomographic bin file, the stage loops through galaxies in each tomographic bin specified by a bin mask and produces the $n(z)$ for each tomographic bin. As pointed out in Appendix A of Rau et al. (2023) this “stacking" is a valid (though not necessarily optimal) ensemble estimate for CDE-based methods, e.g., FlexZBoost. However, this is a ‘naive’ way to obtain the ensemble redshift distribution; as pointed out in Malz & Hogg (2020), it overestimates the photometric redshift uncertainty.
@@ 5.1 Distribution-to-Distribution Metrics
[…]
    where $N$ is the number of galaxies and $F_{N}(z)$ is the empirical cumulative density function for the $N$-th galaxy.
@@ 5.2 Distribution-to-Point Metrics
  • Conditional density loss (CDELoss): we implement the method in Izbicki & Lee (2017a) to compute the mean square error of the difference:

### v3

@@ 1 Introduction
The Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) will obtain deep ugrizyugrizy imaging of $\sim 18,000$ square degrees optimized for extragalactic science during its ten-year survey. It will produce a catalog of 20 billion galaxies down to $i=26.4$, enabling advances in our understanding of numerous extragalactic phenomena and cosmology (Ivezić et al., 2019). Most, if not all, such studies require some notion of the distance or redshift of each galaxy, and in some cases the distribution of redshifts for ensembles of galaxies (see Table 1 of Breivik et al. 2022). Redshift may be measured directly from the absorption and emission lines in a spectrum. However, LSST’s galaxy sample poses a twofold challenge to redshifts measured directly from the absorption and emission lines in a spectrum: the large number of galaxies greatly exceeds the available time on spectroscopic follow-up observing facilities, and the faint sources that comprise the bulk of the sample will be inaccessible to spectroscopic instruments regardless.
In lieu of spectroscopically confirmed redshifts, LSST will yield photometric redshifts (photo-$z$s) derived from broadband photometry, a standard data product of photometric galaxy surveys (e.g., Tanaka et al., 2018; Bilicki et al., 2018; Buchs et al., 2019). Photo-$z$s are subject to a variety of sources of inaccuracy and imprecision that are anticipated to be more severe for LSST’s faint sample, making point estimates and Gaussian uncertainties inappropriate approximations for nearly all extragalactic applications of LSST data. Photo-$z$ probability density functions (PDFs) comprehensively characterize the redshift uncertainty (Tanaka et al., 2018; Bilicki et al., 2018; Buchs et al., 2019) and are thus favored for LSST data. A thorough review of photo-$z$ uncertainty characterization can be found in Newman and Gruen (2022).
@@ 2.3 Data Handle
The tabular data handles (TableHandle) wrap catalog-like data, which can exist as a Numpy dictionary, pandas dataframe, pyarrow table, or Astropy table. The TableHandle interacts with tables_io to handle different types of file formats. More information can be found about tables_io in Appendix B.1. The qp handle (QPHandle) wraps qp ensembles (Malz and Marshall, 2017), which are iterable data structures of 1D PDFs. Additional information about qp can be found in Appendix B.2. The model handle (ModelHandle) wraps the model generated by a particular algorithm in RAIL, informed by some training data. Regardless of their types, the models are stored in pickle files by RAIL.
@@ 2.4 Definitions
Photo-$z$ probability density distribution (PDF) $p(z)$ A photo-$z$ PDF in this paper is referred to the posterior distribution of the redshift of an individual galaxy. This is used interchangeably with $p(z|\mathbf{p})$, where $\mathbf{p}$ is the galaxy’s photometry. The conditions are omitted for notation clarity only.
@@
Photometric data $\mathbf{p}$ Photometric data in this paper is referred to the galaxy catalog information containing photometry (i.e. fluxes or magnitudes in different filters and colours. For some algorithms, this can also include non-photometric information, such as morphology). This can refer to both simulated and real data.
@@
Color $c$ A color is defined as $c=m_{I}-m_{I^{\prime}}$, where $I$ and $I^{\prime}$ are adjacent filters. It is also often referred to as $I-I^{\prime}$. For example, ugu-g or gig-i color.
@@
True redshift PDF $p(z_{t}|\mathbf{p}_{t})$ The conditional PDF of the true redshift, given the true photometry of a galaxy in the simulation, or a normalizing flow trained on a simulation. In RAIL this is provided for galaxies simulated using a probabilistic model for which the joint distribution of redshift and magnitudes from which the galaxies are sampled is known exactly. For example, this is true of galaxies simulated using PZFlow. In Section 5, we also refer to this as $p_{\rm true}(z)$.
@@
Figure 2: The workflow of the RAIL.creation forward modeling subpackage. Input and output data are represented by rectangles, and RAIL stages are represented by ovals. A typical creation pipeline starts with training a creation (‘generative’) model from either a reference catalog (often simulations), such as in the case of PZFlow, or from the template SEDs drawn from a prior, such as in the case of DSPS/FSPS. The empirical or explicit joint distribution of true redshifts with other catalog properties ($\mathbf{p}$) is then computed. A mock truth catalog (‘underlying galaxy catalog’) is then drawn from the creation model. This catalog is then degraded by the degradation stages, such that it mimics a noisy observed catalog (‘biased galaxy sample’).
@@ 3 Creation and Degradation
Mock DESC data are important for systematically testing the performance of various photo-$z$ algorithms. One of the lessons learned from DC1 is that it is desirable for the mock data to include not only true redshifts and LSST photometry (i.e., fluxes in the six LSST bands) but also true posterior PDFs, $p(z_{t}|\mathbf{p}_{t})$, which are unavailable for spectroscopically confirmed data sets as well as traditional simulations. This allows principled comparison with the estimated $p(z|\mathbf{p})$ under non-ideal conditions, such as realistic noise, selection effects, and biases (e.g. see Crafford et al. in prep.). This is critical for the training and validation of photo-$z$ algorithms.
@@ 3.1 Engines
An engine is defined by a pair of stages that are subclasses of each of the following superclasses: RAIL.creation.Modeler makes a model of the $p(z,\mathbf{p})$ joint probability space based on input parameters or data, and RAIL.creation.Creator samples $(z,\mathbf{p})$ from the forward model. Available engines are listed in Table 2.
@@
FSPS Physical rail-fsps Conroy et al. (2009); Conroy and Gunn (2010)
@@ 3.1.1 FSPS (Flexible Stellar Population Synthesis)
FSPS is a RAIL module that creates an interface to the Python bindings of the popular stellar population synthesis (SPS) code FSPS (Flexible Stellar Population Synthesis, Conroy et al. 2009; Conroy and Gunn 2010). FSPS aims at generating realistic galaxy spectral energy distributions (SEDs) by modelling all the components that contribute to the light from a galaxy: stars, gas, dust and AGN. FSPS is widely used both for stellar population inference (Johnson et al., 2021) and for forward modelling of galaxy SEDs (e.g., Alsing et al. 2023; Tortorelli et al. 2024).
@@ 3.2.4 QuantityCut
This degrader provides a trimmed version of the input catalog based on selection cuts applied to the catalog quantities. The user provides the parameter cuts, which is a dictionary with keys being the columns to which the selection is to be applied (e.g., the $i$-band magnitude), and the values being the specific cuts. Two types of values can be provided: a single float number (e.g., 25.3), which is interpreted as a maximum value (i.e., the cut will remove samples with $i>25.3$), and a tuple (e.g., $(17,25.3)$), which is interpreted as a range within which the sample is selected (i.e., the selected sample has $17<i<25.3$). When multiple cuts are applied at the same time, only the intersection of selected samples of each cut will be kept in the output.
@@ 3.2.7 Blending Degrader
Estimator Model ParametersPrior InformationInform EstimatorEstimation ModelTest SetApply EstimatorPhoto-$z$ PDFsOptionally requested photo-$z$ point estimates
@@ 4 Photo-$z$ Estimation
Delight CatInformer, CatEstimator rail-delight Leistedt and Hogg (2017)
@@
FlexZBoost CatInformer, CatEstimator rail-flexzboost Izbicki and Lee (2017b)
@@
TPZ CatInformer, CatEstimator rail-tpz Carrasco Kind and Brunner (2013)
@@
Naive stacking PzInformer, PZSummarizer rail-base Malz and Hogg (2020)
@@ 4.1.3 FlexZBoost
FlexZBoost (Izbicki and Lee, 2017b; Dalmasso et al., 2020) is an algorithm based on conditional density estimate that uses the FlexCode 777available at https://github.com/lee-group-cmu/FlexCode package. The package parameterises the PDF as a linear combination of orthonormal basis functions (a set of unit vectors in the color space that are orthogonal to each other), where the basis function coefficients can be determined by regression. The RAIL implementation uses xgboost (Chen and Guestrin, 2016) to perform the regression. The basis function representation of the photo-z PDF of a galaxy can lead to small-scale residual ‘bumps’. In the course of training the density estimate, an optimal threshold (configuration parameter bump_thresh) below which small-scale features are removed is determined by setting aside a fraction of the training data and minimizing the CDE loss at different threshold values. Additionally, the width of the final PDF is similarly optimized by the inclusion of a ‘sharpening’ parameter that scales the PDF by a power law value $\alpha$. Again a fraction of the training data is set aside and the CDE loss is minimized over a set of $\alpha$ values. The resultant photo-$z$ PDF distributions can be stored as qp.Ensembles either in their native basis function representation, or as a linearly interpolated grid.
@@ 4.1.4 GPz
GPz is an algorithm based on sparse Gaussian Process, which was introduced by Almosallam et al. (2016). The current RAIL implementation of GPz is the preliminary version; that is, it predicts a single Gaussian PDF rather than the more sophisticated multimodal PDFs that are implemented in newer versions of GPz (Stylianou et al., 2022). GPz models both the mean and standard deviation of the Gaussian PDF as a linear combination of basis functions, and learns the parameters for the basis functions via a Gaussian process. The method can make one of several assumptions about the covariance between these basis functions, which are controlled via the configuration parameter gpz_method as outlined in the RAIL documentation.
@@ 4.2.1 BPZ (Bayesian Photometric Redshifts)
BPZ is a template-based estimator developed by Benítez (2000), see that reference and Coe et al. (2006) for more details on the algorithm. Like many template-based codes, it operates by computing synthetic fluxes for an input set of SEDs by integrating the products of the SEDs and the filter bandpass curves for a particular survey. For each SED the algorithm computes a $\chi^{2}$ value using the (scaled) fluxes and the observed data. These $\chi^{2}$ values are converted into likelihoods, and an SED-type-dependent apparent magnitude prior is applied to the likelihoods before they are marginalized over type to produce a final posterior probability distribution. Two additional quantities are included in the ancillary data: tb corresponding to the ‘best-fit SED type’ (evaluated at the mode redshift), and todds, a parameter that gives the fraction of the probability that comes from SED type tb at the mode redshift. Low values of todds may be indicative of a poor fit.
@@ 4.3.1 Delight
Leistedt and Hogg (2017) introduced a novel approach to inferring photometric redshifts which combines some of the strengths of machine learning and template-fitting methods by implicitly constructing flexible template SEDs directly from the spectroscopic training data, called Delight. It is a method for calculating the posterior probability of redshift given a catalog of deep observations acting as a data-driven prior. The catalog can have observations in arbitrary bands and with arbitrary noise; Gaussian processes are used as a principled method to implicitly construct SEDs (capturing the effects of redshifts, bandpasses and noise). The hyperparameters of the Gaussian process can be optimized as a calibration step.
@@ 4.5.3 Naive methods
The NaiveStackSummarizer takes the photo-$z$ PDF from each catalog object and simply average them over the ensemble to produce the ensemble redshift distribution $n(z)$, i.e., $n(z)=\sum_{i}^{N}p_{i}(z)/N$. A set of bootstrap realizations is also produced to estimate the uncertainties on the distributions. The NaiveStackMaskedSummarizer can deal with tomographic bins. Given the tomographic bin file, the stage loops through galaxies in each tomographic bin specified by a bin mask and produces the $n(z)$ for each tomographic bin. As pointed out in Appendix A of Rau et al. (2023) this “stacking" is a valid (though not necessarily optimal) ensemble estimate for CDE-based methods, e.g., FlexZBoost. However, this is a ‘naive’ way to obtain the ensemble redshift distribution; as pointed out in Malz and Hogg (2020), it overestimates the photometric redshift uncertainty.
@@ 5.1 Distribution-to-Distribution Metrics
    ${\rm AD}=N\sum_{i=1}^{N}\int_{0}^{+\infty}\frac{(F_{i}(z)-F(z))^{2}}{F(z)(1-F(z))}\,p(z)dz,$ (10)
    where $N$ is the number of galaxies and $F_{i}(z)$ is the empirical cumulative density function for the $i$-th galaxy.
@@ 5.2 Distribution-to-Point Metrics
  • Conditional density loss (CDELoss): we implement the method in Izbicki and Lee (2017a) to compute the mean square error of the difference: